Grade 11Mechanics

A train travels 5km in the north east direction at the rate of 40km/h and then 5km at the rate of 40km/h in the north west direction. What is its average speed during the whole journey and what is its average velocity.

Kashish
9 Years agoGrade 11

1 Answer

Aman gohel
ApprovedApproved Tutor Answer9 Years ago
Case 1:(For average speed) Av speed=Total d/Total t Time=1/8(for first journey )+1/8(for second journey ) =0.25hrs Total d=5+5=10kmsTherefore,Avr speed=10/0.25 =40Km/hr(Ans)Case 2:(For average velocity) Avr velocity=displacement/total timeNet displacement=√(5^2+5^2)[According to triangle`s law of vector addition]Displacement=√50Therefore, Avr velocity=√50/0.25 =20√2 =28.28Km/hr(Ans)